# SymbolNet: Neural Symbolic Regression with Adaptive Dynamic Pruning

### v2

@@ 4.3 Datasets and experiments
We test our framework on datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$: the Large Hadron Collider (LHC) jet tagging dataset with 16 inputs [41], MNIST with 784 inputs [42], and SVHN with 3072 inputs [43].
@@ 4.3.1 LHC jet tagging
The labels are one-hot encoded. We train $\tt{SymbolNet}$ with five output nodes, each generating an expression corresponding to one of the five jet substructure classes. We consider models with one or two hidden symbolic layers, using unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(\cdot)$, $\exp(-(\cdot)^{2})$, $\sinh(\cdot)$, $\cosh(\cdot)$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct multiple trials, varying the number of operators within the ranges $u\in[2,30]$ for unary operators and $b\in[2,30]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.6,0.99]$, $\alpha_{\text{input}}\in[0.4,0.9]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
@@ 4.3.2 MNIST and binary SVHN
The MNIST dataset consists of grayscale images of handwritten digits ranging from ‘0’ to ‘9’, each with an input dimension of $28\times 28\times 1$. The task is to classify the correct digit in a given image. Fig. 6b shows an example input image for each of the ten classes. The inputs are flattened into a 1D array, denoted $x_{0},...,x_{783}$, and scaled to the range of $[0,1]$. We train $\tt{SymbolNet}$ with ten output nodes, each generating an expression corresponding to one of the ten digit classes. We consider models with one or two hidden symbolic layers, using unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(-(\cdot)^{2})$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct multiple trials by varying the number of operators within the ranges $u\in[2,20]$ for unary operators and $b\in[2,20]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.7,0.999]$, $\alpha_{\text{input}}\in[0.6,0.99]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
Similar to MNIST but more challenging, the SVHN dataset consists of digit images with an input dimension of $32\times 32\times 3$ in RGB format, where the digits are taken from noisy real-world scenes. For simplicity, we focus on binary classification between the digits ‘1’ and ‘7’. Fig. 6c shows some example input images for each of these two classes. The inputs are flattened into a 1D array, denoted $x_{0},...,x_{3071}$,and scaled to the range of $[0,1]$. In this binary setting, we label the digit ‘1’ as 0 and the digit ‘7’ as 1, so $\tt{SymbolNet}$ has one output node and generates one expression per model. We consider a single hidden symbolic layer, with unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(-(\cdot)^{2})$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct multiple trials by varying the number of operators within the ranges $u\in[2,20]$ for unary operatos and $b\in[2,10]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.8,0.999]$, $\alpha_{\text{input}}\in[0.8,0.999]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
@@ 5.1 LHC jet tagging
$g$ $-0.041\text{Multiplicity}\times C_{1}^{\beta=1}+0.53\tanh(0.6\text{Multiplicity}-0.38C_{1}^{\beta=1})+0.24$ 16 0.885
$q$ $0.150.073\text{Multiplicity}\times C_{1}^{\beta=1}-0.38\tanh(0.63m_{\text{mMDT}})+$ 0.15 12 0.827
$t$ $0.240.2\sin(1.2\text{Multiplicity})+0.43\sin(0.49C_{1}^{\beta=2})-0.2\tanh(0.6\text{Multiplicity}-0.38C_{1}^{\beta=1})+$ 0.24 24 0.915
$W$ $-0.099\sin(0.73\text{Multiplicity})+0.84\exp(-46.0(m_{\text{mMDT}}+0.14C_{1}^{\beta=1}+0.27C_{1}^{\beta=2})^{2})+0.044$ 23 0.894
$Z$ $0.43\exp(-6.9(C_{1}^{\beta=2})^{2})$ 8 0.851
@@
Figure 8: Performance of $\tt{SymbolNet}$ on the LHC jet tagging dataset. (a): The Pareto front illustrates the trade-off between accuracy and expression complexity. (b): ROC curves for a compact symbolic model with a mean complexity of 16 and an overall accuracy of 71%, with its expressions tabulated in Tab. 1.
@@ 5.2 MNIST and binary SVHN
Figure 9: Performance of $\tt{SymbolNet}$ on the MNIST and binary SVHN datasets. (a): The Pareto front illustrates the trade-off between accuracy and expression complexity, comparing $\tt{SymbolNet}$ with $\tt{EQL}$ on the MNIST dataset. (b): The ROC AUC is plotted against expression complexity to demonstrate the performance of $\tt{SymbolNet}$ on the binary SVHN dataset. (c): ROC curves for a compact symbolic model with a mean complexity of 90 and an accuracy of 80% on the MNIST dataset (the expressions of this model are listed in Tab. 2). (d): ROC curves for four symbolic models with different complexity values on the binary SVHN dataset (the expression for the model with a complexity of 122 is listed in Tab. 3).
@@ 5.3 FPGA resource utilization and latency
6 Conclusion
We have proposed $\tt{SymbolNet}$, a neural network approach to symbolic regression featuring a novel and dedicated pruning framework designed to generate compact expressions capable of fitting high-dimensional data.
Most existing methods, whether based on genetic programming or deep learning, primarily focus on datasets with input dimensions less than $\mathcal{O}(10)$ and cannot scale efficiently beyond. In genetic programming approaches, the search algorithms create and evolve equations using a combinatorial strategy, which becomes highly inefficient as the equation search space scales exponentially with its building blocks. On the other hand, while deep learning techniques have demonstrated their ability to handle complex datasets in other domains, they have not been extensively explored for symbolic regression in high-dimensional problems.
Our proposed method, equipped with a novel pruning framework dedicated to symbolic regression, attempts to fill this gap. Its effectiveness has been demonstrated on datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$. While there is always a trade-off between prediction accuracy and computational resources, this work provides an option to minimize the latter. It also opens up potential applications of symbolic regression for model compression, enabling the deployment of more compact symbolic models in resource-constrained environments.
However, neural networks currently have certain limitations when used for symbolic regression compared to genetic programming. For example, it is challenging to create a unified framework that can incorporate non-differentiable mathematical operators, such as division and conditionals, without specifically regularizing each of these operators. These operators can introduce singular points in gradient-based optimization, limiting the model’s ability to fit data effectively. Additionally, while genetic programming can generate a batch of models due to its discrete nature, the process is relatively slow. In contrast, neural networks can only generate one model at a time, although training is usually fast due to their continuous nature and the use of gradient descent. These challenges will be addressed in future work.

### v3

@@ 4.3 Datasets and experiments
We test our framework on datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$: the Large Hadron Collider (LHC) jet tagging dataset with 16 inputs [48], MNIST with 784 inputs [49], and SVHN with 3072 inputs [50]. Datasets are split into train, validation, and test sets with ratios of 0.6/0.2/0.2. Models are trained for 200 epochs with a batch size of 1024 using the $\tt{Adam}$ optimizer [51] and a learning rate of 0.0015. As described in Sec. 3.2, trainable thresholds for all pruning types are initialized to zero and are clipped to the range [0,1] when competing with the corresponding fixed threshold target of 1, except for thresholds associated with model weight pruning, which are clipped to [0$\infty$] since the model weights are unbounded. Model weights and biases are initialized with random normal distributions. Typically, one to two symbolic hidden layers with $\mathcal{O}(1-10)$ unary and binary operators are sufficient for most cases. There is flexibility in choosing differentiable operators for activations, with elementary functions like trigonometric and exponential functions being adequate due to the extensive function space the network can represent.
@@ 4.3.1 LHC jet tagging
The labels are one-hot encoded. We train $\tt{SymbolNet}$ with five output nodes, each generating an expression corresponding to one of the five jet substructure classes. We consider models with one or two hidden symbolic layers, using unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(\cdot)$, $\exp(-(\cdot)^{2})$, $\sinh(\cdot)$, $\cosh(\cdot)$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct 40 trials for a grid search of the hyperparameters, varying the number of operators within the ranges $u\in[2,30]$ for unary operators and $b\in[2,30]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.6,0.99]$, $\alpha_{\text{input}}\in[0.4,0.9]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
@@ 4.3.2 MNIST and binary SVHN
The MNIST dataset consists of grayscale images of handwritten digits ranging from ‘0’ to ‘9’, each with an input dimension of $28\times 28\times 1$. The task is to classify the correct digit in a given image. Fig. 6b shows an example input image for each of the ten classes. The inputs are flattened into a 1D array, denoted $x_{0},...,x_{783}$, and scaled to the range of $[0,1]$. We train $\tt{SymbolNet}$ with ten output nodes, each generating an expression corresponding to one of the ten digit classes. We consider models with one or two hidden symbolic layers, using unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(-(\cdot)^{2})$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct 40 trials, for a grid search of hyperparameters, by varying the number of operators within the ranges $u\in[2,20]$ for unary operators and $b\in[2,20]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.7,0.999]$, $\alpha_{\text{input}}\in[0.6,0.99]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
Similar to MNIST but more challenging, the SVHN dataset consists of digit images with an input dimension of $32\times 32\times 3$ in RGB format, where the digits are taken from noisy real-world scenes. These images often include various distractors alongside the digit of interest, making the 10-class classification particularly challenging without additional architectural modifications, such as convolutional layers or other techniques to handle the complexity. Therefore, for simplicity, we focus on binary classification between the digits ‘1’ and ‘7’. Fig. 6c shows some example input images for each of these two classes. The inputs are flattened into a 1D array, denoted $x_{0},...,x_{3071}$,and scaled to the range of $[0,1]$. In this binary setting, we label the digit ‘1’ as 0 and the digit ‘7’ as 1, so $\tt{SymbolNet}$ has one output node and generates one expression per model. We consider a single hidden symbolic layer, with unary operators including $\sin(\cdot)$, $\cos(\cdot)$, $\exp(-(\cdot)^{2})$, and $\tanh(\cdot)$, and binary operators including $+$ and $\times$. To explore a range of expression complexities, we conduct 40 trials, for a grid search of hyperparameters, by varying the number of operators within the ranges $u\in[2,20]$ for unary operatos and $b\in[2,10]$ for binary operators. We also vary the target sparsity ratios within the ranges $\alpha_{\text{weight}}\in[0.8,0.999]$, $\alpha_{\text{input}}\in[0.8,0.999]$, $\alpha_{\text{unary}}\in[0.2,0.5]$, and $\alpha_{\text{binary}}\in[0.2,0.5]$.
@@ 5.1 LHC jet tagging
$g$ $-0.041\text{Multiplicity}\bm{\times}C_{1}^{\beta=1}+0.53\bm{\tanh(}0.6\text{Multiplicity}-0.38C_{1}^{\beta=1}\bm{)}+0.24$ 16 0.885
$q$ $0.150.073\text{Multiplicity}\bm{\times}C_{1}^{\beta=1}-0.38\bm{\tanh(}0.63m_{\text{mMDT}}\bm{)}+$ 0.15 12 0.827
$t$ $0.240.2\bm{\sin(}1.2\text{Multiplicity}\bm{)}+0.43\bm{\sin(}0.49C_{1}^{\beta=2}\bm{)}-0.2\bm{\tanh(}0.6\text{Multiplicity}-0.38C_{1}^{\beta=1}\bm{)}+$ 0.24 24 0.915
$W$ $-0.099\bm{\sin(}0.73\text{Multiplicity}\bm{)}+0.84\bm{\exp(}-46.0\bm{(}m_{\text{mMDT}}+0.14C_{1}^{\beta=1}+0.27C_{1}^{\beta=2}\bm{)}^{\bm{2}}\bm{)}+0.044$ 23 0.894
$Z$ $0.43\bm{\exp(}-6.9\bm{(}C_{1}^{\beta=2}\bm{)}^{\bm{2}}\bm{)}$ 8 0.851
@@
Figure 8: Performance of $\tt{SymbolNet}$ on the LHC jet tagging dataset. (a): The Pareto front, each with models selected from the 40 trials, illustrates the trade-off between accuracy and expression complexity. (b): ROC curves for a compact symbolic model with a mean complexity of 16 and an overall accuracy of 71%, with its expressions tabulated in Tab. 1.
@@ 5.2 MNIST and binary SVHN
Figure 9: Performance of $\tt{SymbolNet}$ on the MNIST and binary SVHN datasets. (a): The Pareto front, each with models selected from the 40 trials, illustrates the trade-off between accuracy and expression complexity, comparing $\tt{SymbolNet}$ with $\tt{EQL}$ on the MNIST dataset. (b): The ROC AUC, with models selected from 40 trials, is plotted against expression complexity to demonstrate the performance of $\tt{SymbolNet}$ on the binary SVHN dataset. (c): ROC curves for a compact symbolic model with a mean complexity of 90 and an accuracy of 80% on the MNIST dataset (the expressions of this model are listed in Tab. 2). (d): ROC curves for four symbolic models with different complexity values on the binary SVHN dataset (the expression for the model with a complexity of 122 is listed in Tab. 3).
@@ 5.3 FPGA resource utilization and latency
6 Discussion
We have introduced $\tt{SymbolNet}$, a neural network-based approach to symbolic regression for model compression, featuring a novel pruning framework designed to generate compact expressions capable of fitting high-dimensional data. Our method aims to address the latency and bandwidth bottlenecks encountered in modern high-energy physics experiments, such as those at the CERN LHC, where datasets are typically high-dimensional and do not have ground-truth equations. By leveraging the efficiency of compact symbolic representations, which can be implemented on custom hardware such as FPGAs, we further extend the applicability of symbolic regression to higher-dimensional data, as commonly found in LHC experiments.
Most existing symbolic regression methods, whether based on genetic programming or deep learning, primarily focus on datasets with input dimensions less than $\mathcal{O}(10)$ and cannot scale efficiently beyond. In genetic programming approaches, the search algorithms create and evolve equations using a combinatorial strategy, which becomes highly inefficient as the equation search space scales exponentially with its building blocks. Even with an external feature selector, the GP-based algorithms can still be impractical for datasets with input dimensions beyond $\mathcal{O}(100)$. On the other hand, while deep learning techniques have demonstrated their ability to handle complex datasets in other domains, they have not been extensively explored for symbolic regression in high-dimensional problems.
Our proposed method, equipped with a novel pruning framework dedicated to symbolic regression, aims to fill this gap. Its effectiveness has been demonstrated on datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$. While there is always a trade-off between prediction accuracy and computational resources, this work provides an option to minimize the latter. Our method is specifically designed to be a model compression technique by leveraging the compact representation of symbolic expressions for low-latency deployment in resource-constrained environments.
However, neural networks currently have certain limitations when used for symbolic regression compared to genetic programming. For example, it is challenging to create a unified framework that can incorporate mathematical operators that are not differentiable everywhere, such as logarithms and division, without specifically regularizing each of these operators. While these operators could expand model expressivity, they can introduce singular points in gradient-based optimization. Potential future extensions include tackling time series or sequence data, integrating domain-specific constraints or prior knowledge, and developing methods to improve the interpretability of generated expressions from high-dimensional datasets. To fully maximize the potential of efficient hardware deployment of symbolic expressions, further work is needed to optimize the expressions while maintaining numerical precision within fixed bit widths, a process known as quantization-aware training [65]. These topics will be addressed in future work.